Alex invested $6000, part at 6% simple interest and part at 4% simple interest for a period of 1 year. How much did he invest at each rate if each account earned the same interest.
Answer by mananth(16949)

let him invest $x at 6%
He will invest 6000-x at 4%
..
interest at6% = 0.06x for 1 year
interest at4% =0.04(6000-x) for 1 year
..
They both should be the same
0.06x=0.04(6000-x)
0.06x=240-0.04x
0.06x+0.04x=240
0.1x = 240
x= $2400 at 6%
& $ 3600 at 4%